you hold a 1 kg basketball 1.5 m above the ground in order to calculate the coefficient of restitution. In this instant, what is the gravitational potential energy of the basketball?

a. 23 J

b. 37 J

c. 72 J

d. 18 J

Using the information from the previous question, if you were to release the basketball, what would be its velocity the moment before it hits the ground?

a. 78 m/s

b. 43 m/s

c. 42 m/s

d. 9.81 m/s
Answer #1

Solution) given

M= 1kg

h = 1.5m

Gravitational potential energy PE = Mgh =1×9.8×1.5=14.7J

PE =14.7 J

NOTE :Kindly check options for above problem

Velocity v=(2gh)^(1/2)=(2×9.8×1.5)^(1/2)= 5.42 m/s

NOTE: Kindly check options